OpenStudy (anonymous):

Can anyone explain in great detail what the Quadractic Formula means and how I can solve Quadratic Equations but for a dummy like me

OpenStudy (anonymous):

I will give medal to first person to explain it to me in a way that I can understand it

OpenStudy (anonymous):

If you wanna know why I am asking these questions it is because I might not get into the schools I am applying to for 6th grade

OpenStudy (anonymous):

even though I got a decent score of about 1900 on the middle school ssat

OpenStudy (anonymous):

I believe I got an 85th percentile on the ssat

OpenStudy (anonymous):

@Mehek14 @pooja195

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

pretty sad that I might not get in

OpenStudy (anonymous):

to the few schools I am applying to

OpenStudy (campbell_st):

The general quadratic formula is used to find the values of x that make the quadratic equation equal zero. There are the points where the parabola cuts of touches the x- axis. All parabola's are symmetric so given a quadratic equation. \[y = ax^2 + bx + c\] the line of symmetry is found using \[x = \frac{-b}{2a}\] so the points that are equi-distance from the line of symmetry can be found using the general quadratic formula. to find the point(s) where the parabola cuts the x-axis set y = 0 so then you have \[0 = ax^2 + bx + c\] the general quadratic formula is \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] you may notice that the line of symmetry equation is part of the quadratic formula.. \[\pm \sqrt{b^2 - 4ac}\] is used to calculate a distance from the line of symmetry on both sides of the line. hope it makes sense.

OpenStudy (anonymous):

so that is why I am learning all this

OpenStudy (anonymous):

can you tell me what you just said but in fewer words and in an easier way to think of it

OpenStudy (danjs):

wow, do you save topic answers or just type all tha tnow.

OpenStudy (campbell_st):

I just typed it off the top of my head @DanJS

OpenStudy (anonymous):

cause I am in middle school and not in 7th grade yet but wanna learn it since I might not get into my dream schools

OpenStudy (danjs):

general quadratic, a,b,c are constant numbers y = ax^2+bx+c when y=0, the graph crosses the x axis 0 = ax^2 + bx + c the formula is just solving that there for x, nice to show it, or people just remember

OpenStudy (anonymous):

is ^ used for exponents?

OpenStudy (danjs):

yes, sorry, good for you learning algebra in 7th grade, dont think i had that till 8th

OpenStudy (anonymous):

also I am not in 7th grade

OpenStudy (anonymous):

I just said my sister learned it in her class for 8th grade or 7th grade so I just wanna try and learn this before I get into her grade to try again at the SSAT and possibly do better

OpenStudy (anonymous):

cause I did get like an 85th percentile on the ssat but no schools look like taking me even though I didn't get any letters yet

OpenStudy (anonymous):

cause I think they just don't like me

OpenStudy (anonymous):

also can you show me on a graph what it would look like

OpenStudy (anonymous):

so I can get a better visual idea of it

OpenStudy (anonymous):

@satellite73

OpenStudy (danjs):

the quadratic form has highest power x^2, these are parabolas, sometimes there might not be any real number solutions to the formula or xintercepts

OpenStudy (anonymous):

I gtg

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

I am back

OpenStudy (anonymous):

@jhonyy9 @triciaal @SolomonZelman @kropot72

OpenStudy (tennis5518):

hope this one help

OpenStudy (danjs):

yeah those are easy videos to follow on that site, start the algebra ones

OpenStudy (anonymous):

I think I saw something that I can understand https://www.youtube.com/watch?v=70dA34KAFWw

OpenStudy (anonymous):

can you check to see if it is legit

OpenStudy (tennis5518):

ok

OpenStudy (anonymous):

is this legit? https://www.youtube.com/watch?v=velArLW85MU

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

I think I got it now

OpenStudy (anonymous):

maybe I can test my skills to see if I know it

OpenStudy (danjs):

i never seen those.. here, play with this site https://www.khanacademy.org/math/algebra-basics make an account on that site , free, start wherever you are comfortable, it is organized into subjects

OpenStudy (tennis5518):

i send him that danJS

OpenStudy (tennis5518):

@DanJS

OpenStudy (anonymous):

could you test me with a problem or two?

OpenStudy (anonymous):

so that I know if I do know it or not

OpenStudy (danjs):

oh right, yeah that is in algebra 1 section, you should be comfortable with all the stuff before that first, then move to those

OpenStudy (danjs):

lines,linear equations. and manipulating equations around, and stuff

OpenStudy (anonymous):

I don't think I understand it

OpenStudy (anonymous):

I really don't understand the quadratic formula and how to solve something like -4x + 2 -4x = 0

OpenStudy (anonymous):

everything I think is just incorrect and I have no idea on how to solve anything

zepdrix (zepdrix):

Would you know how to solve something like this? :) 3x-7+4x=0

OpenStudy (anonymous):

I have no idea @zepdrix

OpenStudy (anonymous):

I really want to learn this

OpenStudy (anonymous):

I need to learn as much algebra as I can and geometry since I might have to go back to my old school

OpenStudy (anonymous):

since no school seems interested in me

OpenStudy (tkhunny):

ax^2 + bx + c = 0 Solve by Completing the Square That's all there is to it.

zepdrix (zepdrix):

Well first, make sure you're comfortable with the idea of `like-terms`. We can combine things together if they are the same in a way. Example: 3x-7 These are different "things" being subtracted, they can't be combined in any meaningful way. But instead if we look at 3x+4x These are just different amounts of "x", whatever x may be. Let's call it an apple. So this example is really 3apples + 4apples, which we know to be 7apples. So therefore, 3x+4x = 7x

zepdrix (zepdrix):

So when we're given a problem like: 3x-7+4x=0 We want to start by combining `like-terms`. We want to combine anything that can be combined. So adding the 4x to the 3x gives us this next step: 7x-7=0

OpenStudy (anonymous):

3x - 7 + 4x = 0 meaning you would do 3x + 4x = 7 making x = 1

zepdrix (zepdrix):

Or add 7 to both sides as you've shown, yes that works just as well :)

zepdrix (zepdrix):

ahh sorry i have to go D:

OpenStudy (anonymous):

so would I do this 3x - 7 + 4x = 0 and then add 7 to both sides giving us 3x+ 4x = 7 and then find out what is a b and c which are a = 3 b = 4 and c = -7 then divide b by 2 then square it so 2 x 2 = 4 so now you have 3x + 4x + 4 which is factorable. You can factor 2 and 2 giving you the solution of 2

OpenStudy (anonymous):

so yea that is the only solution you would have

OpenStudy (anonymous):

If I am right than could you tell me and if I didn't could you show me everything that I did wrong

OpenStudy (danjs):

no, that equation is a line, no x^2 term in there, start https://www.khanacademy.org/math/algebra-basics go thorugh foundations and work through, come here to ask questions for examples

OpenStudy (danjs):

i would recommend, prolly

OpenStudy (kayders1997):

Liam that would work because if you have 3x+4y=7 first the equatoon will have to equal 0 when doing the quadratic formula and you have to have it in the form of x^2+x+c=0

OpenStudy (kayders1997):

*wouldnt

OpenStudy (kayders1997):

If you want to we could do anexample

OpenStudy (kayders1997):

if you have 3x^2-6x=3 so you than you will have to move the 3 to the other side by subtractinf it, so 3x^2-6x-3=0 than you can call a 3 b -6 and c -3 so nuow plug that into the equatoon -b+- b^2-4ac all under the square root divided by 2a

OpenStudy (kayders1997):

So 6 +- 36+36 so (6+-square root 72)/6

OpenStudy (anonymous):

I really don't understand

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@nincompoop @ParthKohli

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

also if you can explain it to me and tell me how to solve it then I will give you a medal and a fan

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (mathstudent55):

A quadratic equation is a 2nd-degree equation usually in only one variable. 2nd degree means that the variable is squared. A typical quadratic equation is a three-term, second degree polynomial (also called a quadratic trinomial) equaling zero. The typical form of a quadratic equation is: \(ax^2 + bx + c = 0\) where x is the variable, and a, b, and c are just numbers. Here are some examples of quadratic equations: \(x^2 + 2x - 5 = 0\) \(2x^2 + 6x - 7 = 0\) \(5v^2 - 2v + 2 = 0\) \(2s^2 + 9 = 0\) \(5w^2 =0\) Notice that all the equations have a term with the variable squared. All quadratic equations are a polynomial of second degree equaling zero. The goal in solving the quadratic equation is to find what values of the variable will make the polynomial equal zero. There are several ways of solving a quadratic equation. One such method is called the complete the square method. The quadratic equation is derived from the complete the square method. Without going through the complete derivation of the quadratic formula, here is the quadratic formula: \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \) In order to use it to solve a quadratic equation, all you need to do is write the quadratic equation is the form \(ax^2 + bx + c = 0\) Then use the numbers a, b, and c in the quadratic formula to find the solutions. Here is an example. Solve using the quadratic formula: \(2x^2 - 9x - 18 = 0\) When you compare this equation with the form \(ax^2 + bx + c = 0\) you see that a = 2, b = -9, and c = -18. Now write the quadratic formula, and then replace a, b, and c with the values from our quadratic equation: \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \) \(x = \dfrac{-(-9) \pm \sqrt{(-9)^2 - 4(2)(-18)}}{2(2)} \) \(x = \dfrac{9 \pm \sqrt{81 + 144}}{4} \) \(x = \dfrac{9 \pm \sqrt{225}}{4} \) \(x = \dfrac{9 \pm 15}{4} \) \(x = \dfrac{9 + 15}{4} \) or \(x = \dfrac{9 - 15}{4} \) \(x = \dfrac{24}{4} \) or \(x = \dfrac{-6}{4} \) \(x = 6 \) or \(x = -\dfrac{3}{2} \) If you test the two solutions in the given equation, you will see that those two values of x will indeed make the value of the polynomial zero.

OpenStudy (whpalmer4):

Think of the quadratic formula as a recipe for giving you the solution that always works. If you can rearrange your formula into the form\[ax^2+bx+c=0\]where \(a \ne 0\) then you can always get the values of \(x\) which make that true by evaluating \[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] There are other ways you can solve a quadratic as well; some are easier to understand or use in some cases. Factoring: if you can factor your quadratic into the form \[k(x-a)(x-b) = 0\]where \(k,a,b\) are constants, then your solutions are simply the values of \(x\) that make \(x-a=0\) and \(x-b=0\) true. For example: \[x^2-3x+2\]can be factored as \[(x-1)(x-2)\]because \[(x-1)(x-2) = x^2-2x-1x+2 = x^2 -3x + 3\]If we set those factors equal to \(0\) and solve for \(x\) we get \[x-2=0, \ x=2\]\[x-1 = 0,\ x=1\] And sure enough, look what happens when we plug those values into our equation: \[x^2-3x+2 = (1)^2-3(1)+2 = 1-3+2 = 0\]\[(2)^2-3(2)+2 = 4-6+2 = 0\]

OpenStudy (whpalmer4):

Completing the square: Another way is what is called "completing the square". Here we have something like \[x^2 + bx + c = 0\] Now imagine we have \[(x+\frac{b}{2})^2 = (x+\frac{b}{2})(x+\frac{b}{2}) = x^2 + x\frac{b}{2} + x\frac{b}{2} + (\frac{b}{2})^2 = x^2 + bx + \frac{b^2}{4}\] That means we could write \(x^2+bx\) as \((x+\frac{b}{2})^2\) as long as we fudge the rest of the equation a bit by simultaneously adding and subtracting \((b/2)^2\). Let me solve the same equation from before: \[x^2-3x+2 = 0\] \[(x^2-3x) + 2 = 0\]We take half of the coefficient of \(x\) (in other words, the number in front of \(x\)), square it, and both add and subtract that number. If we both add and subtract some number, there is no change in the value. If you have $20 in your wallet, and I give you a $10 bill and I take a $10 bill from your wallet, you still have $20. We're going to do the same thing. \[(x^2-3x+(\frac{3}{2})^2) + 2 - (\frac{3}{2})^2 = 0\]\[(x^2-3x+\frac{9}{4})+2-\frac{9}{4} = 0\]The stuff in the parentheses we can rewrite as \((x-\frac{3}{2})^2\) \[(x-\frac{3}{2})^2 + 2 - \frac{9}4=0\]Move the constants to the other side and combine: \[(x-\frac{3}{2})^2 = \frac{9}4-2 = \frac{1}{4}\]Now we take the square root of both sides: \[\pm(x-\frac{3}{2}) = \frac{1}{2}\] I have the \(\pm\) because both the negative and positive square roots give the same result when squared. \[+(x-\frac{3}{2}) = \frac{1}2\]\[x=2\]\[-(x-\frac{3}{2}) =\frac{1}{2}\]\[-x=-1\]\[x=1\] Same solutions as before.

OpenStudy (whpalmer4):

Finally, using the quadratic formula: \[x^2-3x+2=0\]\[ax^2+bx+c=0\]comparing like terms, we can see that \[a=1,\ b=-3,\ c=2\]so our solutions from the quadratic formula are: \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} = \frac{-(-3)\pm\sqrt{(-3)^2-4(1)(2)}}{2(1)} =\frac{3\pm\sqrt{9-8}}{2} \]\[=\frac{3\pm1}{2} = \frac{4}2,\frac{2}2=2,1 \] The quadratic formula is especially useful for real-world problems where the coefficients are decimals, maybe something like \[-0.254x^2 +5.9730x - 23.990141 = 0\]where attempting to factor or complete the square would be a big headache, even if you turned those decimals into fractions, such as \[-\frac{254}{1000}x^2+\frac{5973}{1000}-\frac{23990141}{1000000}=0\]

OpenStudy (anonymous):

I have no idea still and wish to learn

OpenStudy (jdoe0001):

hehhe

OpenStudy (jdoe0001):

hmm have you covered quadratic equations yet? I mean... say... can you throw me one, any one, make any up

OpenStudy (anonymous):

no but wish to since I am applying to prestigious east coast schools so I am wanting to learn how to do this when I am going into 6th grade

OpenStudy (anonymous):

and also I have the option to skip grades in math

OpenStudy (anonymous):

so I am trying to learn this one by one

OpenStudy (jdoe0001):

well... if you hmmm haven't covered quadratics yet this will be a good time to go through them for that is what the quadratric formula is for so. you want to cover quadratics firstly

OpenStudy (anonymous):

found you!

OpenStudy (anonymous):

I know the quadratic formula and how to solve a problem it is just I know some but not others

OpenStudy (anonymous):

what were saying about medals

OpenStudy (anonymous):

oh that you press best response to give somebody one.

OpenStudy (anonymous):

I hope you understand what I mean